• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Cummings

    Log entry time 10:07:01 on April21,2012

    Entry number 377091

    This entry is a followup to: 377067

    keyword=Moving to production - looks like beam is scraping beam pipe

    After the down to reset the left septum, Pengjia and Kalyan worked on centering the beam, but it seems the beam is scraping the beam pipe somewhere (see entry 377067). HRS spot still looks uniform.

    MCC is planning to stop beam at ~11 (or 12?), so in the interest of time we will polarize the target and take some production data. After the next down we will check to see if the beam is still scraping.